OpenStudy (anonymous):

Deanna wants to borrow $8000. Two banks offer a 4-year loan term. One loan has a 5% interest rate, and the other has a 6% interest rate. How much more interest will Deanna pay if she accepts the loan at 6% for the 4-year term?

OpenStudy (arcticfoxzx87):

need some help

OpenStudy (anonymous):

Yeah.

OpenStudy (arcticfoxzx87):

I think it is 0.75 percent

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

It seems like the question was asking for the amount of interest instead of a percentage

OpenStudy (arcticfoxzx87):

oh ok

OpenStudy (arcticfoxzx87):

let me see again

OpenStudy (snowsurf):

So they are asking for the amount that you will be paying in interest if you choose the bank that has 6% interest rate. We can use the interest rate formula. Let's call bank A at 5% and bank B at 6%. \[I_A=Prt\] and \[I_B=Prt\] in order to how much more you will be paying at 6% we need to subtract the bank A from bank B \[I_B-I_A=I\] Where I is the interest you will be paying more. \[I_B-I_A=P(rt_B-rt_A)\] P is the amount that is being borrowed, r is the interest rate and t is the time in years. \[I_B-I_A=8000*(0.06(4)-0.05(4)) \] \[I_B-I_A=8000*(0.24-0.04) \] \[I_B-I_A=8000*(0.04) \] \[I_B-I_A= 320\] Since we let \[I_B-I_A=I\] We can say that \[I=320\] So at bank B you will be paying $320 if you were to barrow $ 8000.